# SustainDiffusion: Optimising the Social and Environmental Sustainability of Stable Diffusion Models

### v1

@@ Abstract.
Background: Text-to-image generation models are widely used across numerous domains. Among these models, Stable Diffusion (SD) – an open-source text-to-image generation model – has become the most popular, producing over 12 billion images annually. However, the widespread use of these models raises concerns regarding their social and environmental sustainability. Aims: To reduce the harm that SD models may have on society and the environment, we introduce SustainDiffusion, a search-based approach designed to enhance the social and environmental sustainability of SD models. Method: SustainDiffusion searches the optimal combination of hyperparameters and prompt structures that can reduce gender and ethnic bias in generated images while also lowering the energy consumption required for image generation. Importantly, SustainDiffusion maintains image quality comparable to that of the original SD model. Results: We conduct a comprehensive empirical evaluation of SustainDiffusion, testing it against six different baselines using 56 different prompts. Our results demonstrate that SustainDiffusion can reduce gender bias in SD3 by 68%, ethnic bias by 59%, and energy consumption (calculated as the sum of CPU and GPU energy) by 48%. Additionally, the outcomes produced by SustainDiffusion are consistent across multiple runs and can be generalised to various prompts. Conclusions: With SustainDiffusion, we demonstrate how enhancing the social and environmental sustainability of text-to-image generation models is possible without fine-tuning or changing the model’s architecture.
Image Generation, Stable Diffusion, Bias Mitigation, Energy Efficiency, SBSE
ccs: Software and its engineering Extra-functional propertiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Search methodologies
@@ 1. Introduction
Image generation models are nowadays employed in several domains, including advertisements, education, or web content generation (Ko et al., 2023). To give an idea of the magnitude of the adoption of these models, a recent survey reported that more than 15 billion images have been generated by text-to-image models between 2022 and 2023 (Valyaeva, 2023). Stable Diffusion (SD), an open-source text-to-image generation model (Ho and Salimans, 2022), was reported as the one mainly employed by end-users, with more than 12 billion images generated through its use in those two years (Valyaeva, 2023). Given the extensive adoption of this model, ensuring its sustainability is paramount.
@@
Software sustainability is defined as ”the preservation of the long-term and beneficial use of software and its appropriate evolution in a context that continuously changes” (Lago et al., 2015). Among the different dimensions in which software sustainability is addressed (Lago et al., 2015), the social and environmental dimensions gained considerable relevance in the Software Engineering (SE) community in recent years (Cruz et al., 2025; Verdecchia et al., 2021; Hort and Sarro, 2021; Brun and Meliou, 2018). Social sustainability is related to mitigating the harm that software systems may have on society in terms of discrimination and bias (Lago et al., 2015). In this matter, it has been shown that SD models expose a significant gender and ethnicity bias when asked to generate images for specific occupations or domains (Bianchi et al., 2023; Naik and Nushi, 2023). Focusing on the SE domain, previous work highlighted how SD models have a significant bias towards Male and White/Asian figures when asked to generate images of a software engineer (Fadahunsi et al., 2025; Bano et al., 2025; Sami et al., 2023). Thus, adopting SD models without specific awareness could amplify the already existing perception of bias in the SE community (d’Aloisio et al., 2023; D’Angelo et al., 2024; Sesari et al., 2024b; Rodríguez-Pérez et al., 2021; Sesari et al., 2024a; Cutrupi et al., 2023, 2024). Environmental sustainability refers instead to the impact that software systems can have on the environment in terms of energy and resource consumption (Lago et al., 2015; Menzies and Johnson, 2024). In this context, it has been shown that generative models consume significant energy (Georgiou et al., 2022; Nguyen et al., 2025; Cheung et al., 2025). For example, a recent research revealed that generating a single image from an image generation model such as SD could use up to 4.08 Wh, which is roughly the energy required to charge a phone up to 40% (Bertazzini et al., 2025).
@@
The different dimensions of sustainability are generally in conflict with each other, and addressing one dimension may lead to worse results in others (Lago et al., 2015). For instance, to increase the environmental sustainability of SD models, data scientists may derive compressed versions of these models (d’Aloisio et al., 2024b). However, compressing these models may exacerbate the gender and ethnicity bias exposed by them. Similarly, a possible solution to increase the social sustainability of SD models could be fine-tuning them on a more diverse dataset. However, it may not solve, or even reduce, their environmental sustainability. Additionally, not all users may have the resources needed to compress or fine-tune an SD model. Therefore, optimising the social and environmental sustainability of SD models can be thought of as a multi-objective problem, where the optimal solutions provide the best trade-off between these objectives.
@@
In contexts like this one, search-based software engineering (SBSE) has been proven to be a widely effective and efficient approach to tackle such tasks (Harman et al., 2008; Cruz et al., 2025; Sarro, 2023a, b). In SBSE, an SE problem is formulated as a search task, and the optimal solutions are identified inside a space of possible solutions through one or more fitness functions that guide this search.
@@
We observe that the social and environmental sustainability of these models can be reduced to a software configuration problem, which makes achieving sustainable models a search-based software optimisation task (Sarro, 2023b).
@@
In this paper we propose SustainDiffusion, a search-based approach that is capable of optimising the social and environmental sustainability of SD models without impacting the quality of the images generated. Specifically, SustainDiffusion searches for the optimal configuration of SD hyperparameters and prompt structures capable of reducing gender and ethnicity bias in the generated images, as well as the energy required to generate them, while keeping, or even improving, the original image quality.
@@
Figure 1 shows an example of the application of SustainDiffusion. Figure 1(a) shows a set of images generated by Stable Diffusion 3 (SD3) with its default hyperparameters for the prompt ”Photo portrait of a Software Engineer who writes documentations”. As also highlighted in previous studies (Fadahunsi et al., 2025; Sami et al., 2023), this prompt structure exposes a significant gender and ethnic bias in the images generated by the default SD3 model. On the other hand, Figure 1(b) shows images generated from the same input prompt by an SD3 model optimised by SustainDiffusion. We observe how SustainDiffusion allows a fairer gender and ethnic distribution in the images generated while keeping an image quality comparable to the original SD3 model. Additionally, as we show in Section 5, SustainDiffusion significantly reduces the amount of energy required to generate a single image. Specifically, we perform an extensive empirical evaluation and show how SustainDiffusion can reduce the gender bias of SD3 models by 68% and the ethnicity bias by 59%. Moreover, we show how SustainDiffusion can reduce the energy consumption, computed as the sum of CPU and GPU energy (Schmidt et al., 2020), of the default SD3 model by 48%.
@@
  • A replication package providing a Python implementation of SustainDiffusion and the results of our evaluation (rep, 2025).
@@ 2. Related Work
Text-to-image generation models like DALL-E and SD have been shown to expose harmful social biases. Studies by Bianchi et al. (Bianchi et al., 2023) and Naik et al. (Naik and Nushi, 2023) show that even neutral prompts result in biased outputs, while Sun et al. (Sun et al., 2024) and Luccioni et al. (Luccioni et al., 2023) document underrepresentation of women and minority groups in images generated for specific occupations.
Recently, studies have focused on the bias exposed by image generation models to the SE domain, motivated by previous studies that highlight the already existing discrimination in this field (d’Aloisio et al., 2023; D’Angelo et al., 2024; Sesari et al., 2024b; Rodríguez-Pérez et al., 2021; Sesari et al., 2024a). Sami et al. found that prompts related to SE roles returned heavily male-skewed images in DALL-E 2 (Sami et al., 2023). Fadahunsi et al. extended the study of Sami et al. on SD models, highlighting how these models expose a significant bias towards Male and White/Asian figures when prompted to generate a software engineer (Fadahunsi et al., 2025). Finally, Bano et al. performed a similar study on GPT-4 and Microsoft’s Copilot models, highlighting how these models also expose a significant bias towards Male and Caucasian figures when generating a software engineer (Bano et al., 2025).
To mitigate such biases, prompt engineering and diversity augmentation techniques have been proposed. Esposito et al. developed Diversity Fine-Tuned prompts to improve fairness metrics across sensitive attributes (Esposito et al., 2023). Bansal et al. (Bansal et al., 2022) and Sakurai et al. (Sakurai and Sato, 2025) further explored the role of phrasing and distributional rebalance in reducing bias.
@@
With the growing environmental impact of AI models, achieving energy efficiency has become a critical research goal under the ”Green AI” initiative (Schwartz et al., 2020). Improving energy efficiency not only reduces carbon footprints but also lowers operational costs, promoting sustainability in AI deployment. In generative AI, achieving energy efficiency is particularly challenging due to the intensive computational requirements of diffusion models. In this matter, Georgiou et al.(Georgiou et al., 2022) and Tu et al. (Tu et al., 2023) have evaluated the energy consumption across deep learning models and hardware configurations. Additionally, Bertazzini et al. have recently analysed the energy consumption of image generation models highlighting their high energy consumption (Bertazzini et al., 2025).
Research into energy efficiency in AI-based systems typically follows two primary approaches: hardware-based and software-based optimisations. Hardware-based methods involve quantization, model pruning, and sparsity augmentation, which reduce the computational load by modifying model architectures or utilising specialised hardware (e.g., TPUs, edge devices) (Choi et al., 2024; Lai et al., 2024; Zhou et al., 2023; Mazurek et al., 2024). While effective, these methods often require extensive retraining or specific hardware configurations, limiting their generalisability. Additionally, recent research has shown how quantization may even increase the energy consumption of generative models (Bertazzini et al., 2025).
On the software side, strategies such as hyperparameter tuning, algorithmic efficiency improvements, and energy-aware scheduling offer flexible optimisation opportunities without altering model architectures (Bambagini et al., 2016). Specific to SD models, Gong et al. proposed an approach that applies search-based hyperparameter tuning on SD models to reduce image generation time while keeping high image quality (Gong et al., 2024).
In this paper, we follow a similar approach by searching for the best hyperparameter configuration and prompt structure to reduce both the energy consumption and bias of SD models.
@@ 3. Methodology
Figure 2. Overview of SustainDiffusion
Figure 2 reports an overview of the SustainDiffusion approach. The first step in the process is the initialisation of a randomly generated population with no repeated individuals. Each individual in the population is represented as a set of model hyperparameters and specific prompt keywords with associated weights (see Section 3.1). Hyperparameters are directly set into the SD model, while keywords and weights are used to modify the input prompt through prompt engineering and prompt weighting (Berger et al., 2023). Next, for each individual, the SD model generates 20 different images using the same input prompt. Following this generation process, the energy consumption of the model, as well as the gender and ethnic bias and the quality of the generated images are computed. These values are used to evaluate the fitness of each individual (see Section 3.3). If the termination criteria are not met, variation in the population is applied using crossover, mutation, and selection operators and the new individuals are evaluated (see Section 3.2). When the termination criteria are met, the individuals achieving the best trade-off between image quality, fairness, and energy efficiency are returned.
In the following section, we describe in detail the proposed approach by discussing the key ingredients of each search-based approach: the individual representation, the search operators, and the adopted fitness functions (Sarro et al., 2016; Moussa et al., 2022).
@@ 3.1. Representation
Number of Inference Steps $\{25;80\}$ with a step of $1$
Positive Keywords $\{0;20\}$ with a step of $1$
Negative Keywords $\{0;25\}$ with a step of $1$
Prompt Weights $\{0;5\}$ with a step of $1$
@@ 3.1.1. Hyperparameters
Our choice of hyperparameters and range values to be explored is driven by previous work employing search-based approaches to improve the quality of images generated by SD models (Berger et al., 2023; Gong et al., 2024). Specifically, we consider guidance scale and interference steps. Guidance scale defines how close a generated image is to the textual input prompt (Saharia et al., 2022). High guidance scale values cause the model to generate images closer to the input prompt, usually at the expense of lower image quality (Ho and Salimans, 2022). The default value is $7.0$. In this study, we explore the range $\{0:20\}$ with a step of $0.1$. Number of inference steps represents the number of denoising steps to generate the image from the model’s latent space (Ho and Salimans, 2022). Higher values can enable the model to generate higher quality images, but at a higher inference time and energy consumption. The default value is $50$; we explore the range $\{25:80\}$ with $1$ as step.
@@ 3.1.2. Prompt Keywords and Weights
Following previous work (Berger et al., 2023; Gong et al., 2024; Iglesias et al., 2025), we employ prompt engineering techniques to further enhance the quality of the generated images. Specifically, we use a selection of positive and negative keywords that are randomly chosen for each individual. Positive keywords are added to the input prompt, while negative keywords form a negative prompt containing concepts that the model should avoid when generating an image. Positive keywords are terms that have been shown to improve the quality of the generated images (e.g., photograph, photoreal, award-winning) (PromptHero, 2023; Iglesias et al., 2025). Conversely, negative keywords are words that tend to diminish the quality of an image (e.g., illustration, painting, drawing) (PromptHero, 2023; Iglesias et al., 2025). To address gender and ethnic bias, we also include keywords that have been shown to influence the representation of gender and ethnicity in the generated images (e.g., ambitious, intelligent, supportive) (Luccioni et al., 2023). Since we aim for a balanced distribution of gender and ethnicity in the images, these keywords are considered for both positive and negative sets. However, we implement a check to ensure that a keyword will not appear in both the positive and negative sets for any individual. In this study, we explore a range of randomly chosen keywords to include that go from 0 to the maximum number of keywords in the positive (20) and negative (25) sets. [^1] The full set of keywords is available in our replication package (rep, 2025)
Finally, in addition to prompt engineering, we utilise prompt weighting to enhance image quality and further reduce gender and ethnic bias. Prompt weighting is a technique that involves assigning weights to specific words in the prompt, thereby increasing their relevance in the model (HuggingFace, 2024). Technically, weights are indicated by appending a number of ‘$+$’ signs after a word in the prompt. The more ‘$+$’ signs after a word, the greater its relevance to the model. In SustainDiffusion, we implemented prompt weighting as an integer assigned to each individual, indicating the weight to apply to each positive and negative keyword. While we could have assigned different weights to various keywords, we opted for uniform weights across all keywords in the positive and negative sets to avoid increasing our search space further. In this study, we explore a range of weights from $\{0:5\}$, with increments of 1.

### v2

@@ Abstract.
Background: Text-to-image generation models are widely used across numerous domains. Among these models, Stable Diffusion (SD) – an open-source text-to-image generation model – has become the most popular, producing over 12 billion images annually. However, the widespread use of these models raises concerns regarding their social and environmental sustainability. Aims: To reduce the harm that SD models may have on society and the environment, we introduce SustainDiffusion, a search-based approach designed to enhance the social and environmental sustainability of SD models. Method: SustainDiffusion searches the optimal combination of hyperparameters and prompt structures that can reduce gender and ethnic bias in generated images while also lowering the energy consumption required for image generation. Importantly, SustainDiffusion maintains image quality comparable to that of the original SD model. Results: We conduct a comprehensive empirical evaluation of SustainDiffusion, testing it against six different baselines using 56 different prompts. Our results demonstrate that SustainDiffusion can reduce gender bias in SD3 by 68%, ethnic bias by 59%, and energy consumption (calculated as the sum of CPU and GPU energy) by 48%. Additionally, the outcomes produced by SustainDiffusion are consistent across multiple runs and can be generalised to various prompts depicting human-like figures in different tasks. Conclusions: With SustainDiffusion, we demonstrate how enhancing the social and environmental sustainability of text-to-image generation models is possible without fine-tuning or changing the model’s architecture.
Image Generation Model, Stable Diffusion, Bias Mitigation, Energy Efficiency, Search-Based Software Engineering, Multi-Objective Optimisation, Empirical Study
copyright: acmlicensedjournalyear: 2026conference: 48th IEEE/ACM International Conference on Software Engineering; April 12 – 17, 2026; Rio de Janeiro, Braziljournalyear: 2026conference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazilccs: Software and its engineering Extra-functional propertiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Search methodologies
@@ 1. Introduction
Image generation models are nowadays employed in several domains, including advertisements, education, or web content generation (10.1145/3581641.3584078). To give an idea of the magnitude of the adoption of these models, a recent survey reported that more than 15 billion images have been generated by text-to-image models between 2022 and 2023 (noauthor_ai_2023). Stable Diffusion (SD), an open-source text-to-image generation model (ho2022classifierfreediffusionguidance), was reported as the one mainly employed by end-users, with more than 12 billion images generated through its use in those two years (noauthor_ai_2023). Given the extensive adoption of this model, ensuring its sustainability is paramount.
@@
Software sustainability is defined as ”the preservation of the long-term and beneficial use of software and its appropriate evolution in a context that continuously changes” (lago2015framing). Among the different dimensions in which software sustainability is addressed (lago2015framing), the social and environmental dimensions gained considerable relevance in the Software Engineering (SE) community in recent years (Cruz25CommunicationsACM; verdecchia2021green; hort2021did; 10.1145/3236024.3264838). Social sustainability is related to mitigating the harm that software systems may have on society in terms of discrimination and bias (lago2015framing). In this matter, it has been shown that SD models expose a significant gender and ethnicity bias when asked to generate images for specific occupations or domains (bianchi2023easily; naik2023social). Focusing on the SE domain, previous work highlighted how SD models have a significant bias towards Male and White/Asian figures when asked to generate images of a software engineer (fadahunsi_how_2025; bano2025doessoftwareengineerlook; sami_case_2023). Thus, adopting SD models without specific awareness could amplify the already existing perception of bias in the SE community (d2023data; d2024uncovering; 05a55879837848539d04ba48ec33b3ad; rodriguez2021perceived; sesari2024givingmajorsatisfactionfairness; cutrupi2023draw; cutrupi2024draw). Environmental sustainability refers instead to the impact that software systems can have on the environment in terms of energy and resource consumption (lago2015framing; SarroIEEEInterview2024). In this context, it has been shown that generative models consume significant energy (georgiou2022green; nguyen2025device; cheung2025comparative). For example, a recent study revealed that generating a single image from an image generation model such as SD could use up to 4.08 Wh, which is roughly the energy required to charge a phone up to 40% (bertazzini2025hidden).
@@
The different dimensions of sustainability are generally in conflict with each other, and addressing one dimension may lead to worse results in others (lago2015framing). For instance, to increase the environmental sustainability of SD models, data scientists may derive compressed versions of these models (d2024compression). However, compressing these models may exacerbate the gender and ethnicity bias exposed by them. Similarly, a possible solution to increase the social sustainability of SD models could be fine-tuning them on a more diverse dataset. However, it may not solve, or even reduce, their environmental sustainability. Additionally, not all users may have the resources needed to compress or fine-tune an SD model. Therefore, optimising the social and environmental sustainability of SD models can be thought of as a multi-objective problem, where the optimal solutions provide the best trade-off between these objectives.
@@
In contexts like this one, search-based software engineering (SBSE) has been proven to be a widely effective and efficient approach to tackle such tasks (harman2008search; Cruz25CommunicationsACM; SarroICPC; SarroRE; fairrf). In SBSE, an SE problem is formulated as a search task, and the optimal solutions are identified inside a space of possible solutions through one or more fitness functions that guide this search.
@@
We observe that the social and environmental sustainability of these models can be reduced to a software configuration problem, which makes achieving sustainable models a search-based software optimisation task (SarroRE).
@@
In this paper, we propose SustainDiffusion, a search-based approach that is capable of optimising the social and environmental sustainability of SD models without impacting the quality of the images generated. Specifically, SustainDiffusion searches for the optimal configuration of SD hyperparameters and prompt structures capable of reducing gender and ethnicity bias in the generated images, as well as the energy required to generate them, while keeping, or even improving, the original image quality.
@@
Figure 1 shows an example of the application of SustainDiffusion. Figure 1(a) shows a set of images generated by Stable Diffusion 3 (SD3) with its default hyperparameters for the prompt ”Photo portrait of a Software Engineer who writes documentations”. As also highlighted in previous studies (fadahunsi_how_2025; sami_case_2023), this prompt structure exposes a significant gender and ethnic bias in the images generated by the default SD3 model. On the other hand, Figure 1(b) shows images generated from the same input prompt by an SD3 model optimised by SustainDiffusion. We observe how SustainDiffusion allows a fairer gender and ethnic distribution in the images generated while keeping an image quality comparable to the original SD3 model. Additionally, as we show in Section 5, SustainDiffusion significantly reduces the amount of energy required to generate a single image. Specifically, we perform an extensive empirical evaluation and show how SustainDiffusion can reduce the gender bias of SD3 models by 68% and the ethnicity bias by 59%. Moreover, we show how SustainDiffusion can reduce the energy consumption, computed as the sum of CPU and GPU energy (codecarbon), of the default SD3 model by 48%.
@@
  • A replication package providing a Python implementation of SustainDiffusion and the results of our evaluation (repl_package; d_aloisio_2025_17831454).
@@ 2. Related Work
Text-to-image generation models like DALL-E and SD have been shown to expose harmful social biases. Studies by Bianchi et al. (bianchi2023easily) and Naik et al. (naik2023social) show that even neutral prompts result in biased outputs, while Sun et al. (sun2024smiling) and Luccioni et al. (luccioni_stable_2023) document underrepresentation of women and minority groups in images generated for specific occupations.
Recently, studies have focused on the bias exposed by image generation models to the SE domain, motivated by previous studies that highlight the already existing discrimination in this field (d2023data; d2024uncovering; 05a55879837848539d04ba48ec33b3ad; rodriguez2021perceived; sesari2024givingmajorsatisfactionfairness). Sami et al. found that prompts related to SE roles returned heavily male-skewed images in DALL-E 2 (sami_case_2023). Fadahunsi et al. extended the study of Sami et al. on SD models, highlighting how these models expose a significant bias towards Male and White/Asian figures when prompted to generate a software engineer (fadahunsi_how_2025). Finally, Bano et al. performed a similar study on GPT-4 and Microsoft’s Copilot models, highlighting how these models also expose a significant bias towards Male and Caucasian figures when generating a software engineer (bano2025doessoftwareengineerlook).
To mitigate such biases, prompt engineering and diversity augmentation techniques have been proposed. Esposito et al. developed Diversity Fine-Tuned prompts to improve fairness metrics across sensitive attributes (esposito2023mitigating). Bansal et al. (bansal2022ethical), Sakurai et al. (sakurai2025fairt2i), and Friedrich et al. (friedrich2023fair) further explored the role of phrasing and distributional rebalance in reducing bias. Finally, Jiang et al. (10.1145/3664647.3680748) and Yu et al. (yu_bimodal_2025) investigated the topic of bias mitigation at the embedding level by modifying both the prompt and image embeddings.
@@
With the growing environmental impact of AI models, achieving energy efficiency has become a critical research goal under the ”Green AI” initiative (schwartz2020green). Improving energy efficiency not only reduces carbon footprints but also lowers operational costs, promoting sustainability in AI deployment. In generative AI, achieving energy efficiency is particularly challenging due to the intensive computational requirements of generative models. In this matter, Georgiou et al. (georgiou2022green) and Tu et al. (tu2023unveiling) have evaluated the energy consumption across deep learning models and hardware configurations. Concerning diffusion models, Rombach et al. highlight in the original paper proposing the Diffuser architecture how there is a trade-off between image quality and inference time (Rombach_2022_CVPR). Additionally, Bertazzini et al. have recently analysed the energy consumption of image generation models, highlighting their high energy consumption (bertazzini2025hidden).
Research into energy efficiency in AI-based systems typically follows different approach categories (jarvenpaa2024synthesis), such as input data optimisation (verdecchia2022data), designing optimised algorithms (shanbhag2022towards), model optimisation (gong2024greenstableyolooptimizinginferencetime), approaches for energy-aware training (kim2022tradeoff), methods to optimise the deployment of AI models (kim2022tradeoff), and energy-aware model management (poenaru2023retrain).
Our approach falls within the model optimisation category, where we search for the optimal inference hyperparameter configuration and prompt structure to minimise both the energy consumption and bias of SD models. Other techniques in this category may involve model quantisation, pruning and sparsity augmentation, which reduce the computational load by modifying the internal model architectures. While effective, these methods often require extensive retraining or specific hardware configurations, limiting their generalisability. Additionally, recent research has shown how quantisation may even increase the energy consumption of generative models (bertazzini2025hidden). Instead, our proposed approach treats SD models as a black box, not requiring changing the internal model architecture or fine-tuning the model.
Finally, specific to SD models, Gong et al. proposed an approach that applies search-based hyperparameter tuning on SD models to reduce image generation time while keeping high image quality (gong2024greenstableyolooptimizinginferencetime). However, unlike our approach, they do not take into account energy consumption, as well as gender and ethnic bias.
@@ 3. Methodology
Figure 2 reports an overview of the SustainDiffusion approach. The first step in the process is the initialisation of a randomly generated population with no repeated individuals. Each individual in the population is represented as a set of model hyperparameters and specific prompt keywords with associated weights (see Section 3.1). Hyperparameters are directly set into the SD model, while keywords and weights are used to modify the input prompt through prompt engineering and prompt weighting (berger2023stableyolo). Next, for each individual, the SD model generates 20 different images using the same input prompt. Following this generation process, the energy consumption of the model, as well as the gender and ethnic bias and the quality of the generated images are computed. These values are used to evaluate the fitness of each individual (see Section 3.3). If the termination criteria are not met, variation in the population is applied using crossover, mutation, and selection operators and the new individuals are evaluated (see Section 3.2). When the termination criteria are met, the individuals achieving the best trade-off between image quality, fairness, and energy efficiency are returned.
In the following section, we describe in detail the proposed approach by discussing the key ingredients of each search-based approach: the individual representation, the search operators, and the adopted fitness functions (sarro_multi-objective_2016; moussa_meg_2022).
@@ 3.1. Representation
Number of Inference Steps $\{25;80\}$ with a step of 11
Positive Keywords $\{0;20\}$ with a step of 11
Negative Keywords $\{0;25\}$ with a step of 11
Prompt Weights $\{0;5\}$ with a step of 11
@@ 3.1.1. Hyperparameters
Our choice of hyperparameters and range values to be explored is driven by previous work employing search-based approaches to improve the quality of images generated by SD models (berger2023stableyolo; gong2024greenstableyolooptimizinginferencetime). Specifically, we consider guidance scale and inference steps. Guidance scale defines how close a generated image is to the textual input prompt (saharia2022photorealistic). High guidance scale values cause the model to generate images closer to the input prompt, usually at the expense of lower image quality (ho2022classifierfreediffusionguidance). The default value is $7.0$. In this study, we explore the range $\{0:20\}$ with a step of $0.1$. Number of inference steps represents the number of denoising steps to generate the image from the model’s latent space (ho2022classifierfreediffusionguidance). Higher values can enable the model to generate higher quality images, but at a higher inference time and energy consumption. The default value is $50$; we explore the range $\{25:80\}$ with $1$ as step.
@@ 3.1.2. Prompt Keywords and Weights
Following previous work (berger2023stableyolo; gong2024greenstableyolooptimizinginferencetime; Iglesias25), we employ prompt engineering techniques to further enhance the quality of the generated images. Specifically, we use a selection of positive and negative keywords that are randomly chosen for each individual. Positive keywords are added to the input prompt, while negative keywords form a negative prompt containing concepts that the model should avoid when generating an image. Positive keywords are terms that have been shown to improve the quality of the generated images (e.g., photograph, photoreal, award-winning) (prompther; Iglesias25). Conversely, negative keywords are words that tend to diminish the quality of an image (e.g., illustration, painting, drawing) (prompther; Iglesias25). To address gender and ethnic bias, we also include keywords that have been shown to influence the representation of gender and ethnicity in the generated images (e.g., ambitious, intelligent, supportive) (luccioni_stable_2023). Since we aim for a balanced distribution of gender and ethnicity in the images, these keywords are considered for both positive and negative sets. However, we implement a check to ensure that a keyword will not appear in both the positive and negative sets for any individual. In this study, we explore a range of randomly chosen keywords to include that go from 0 to the maximum number of keywords in the positive (20) and negative (25) sets. [^1] The full set of keywords is available in our replication package (repl_package)
Finally, in addition to prompt engineering, we utilise prompt weighting to enhance image quality and further reduce gender and ethnic bias. Prompt weighting is a technique that involves assigning weights to specific words in the prompt, thereby increasing their relevance in the model (prompt_weighting). Technically, weights are indicated by appending a number of ‘++’ signs after a word in the prompt. The more ‘++’ signs after a word, the greater its relevance to the model. In SustainDiffusion, we implemented prompt weighting as an integer assigned to each individual, indicating the weight to apply to each positive and negative keyword. While we could have assigned different weights to various keywords, we opted for uniform weights across all keywords in the positive and negative sets to avoid increasing our search space further. In this study, we explore a range of weights from $\{0:5\}$, with increments of 1.